OpenStudy (anonymous):

Part 1 [3 points]: Find the solution to the equation 64^(x - 4) = 4^2x. Part 2 [4 points]: Use complete sentences to explain the method used to solve this equation.

OpenStudy (campbell_st):

64 = 4^3 so \[64^{x - 4} = 4^{3(x - 4)}\] equating powers since the same base 3(x -4) = 2x 3x - 12 = 2x x = 12

OpenStudy (dumbcow):

take log of both sides use log property: log( x^n) = n*log(x) \[(x-4) \log 64 = 2x \log 4\]

OpenStudy (dumbcow):

or that :{

OpenStudy (anonymous):

Thanks:)